OpenStudy (anonymous):

Karla is given the equation 10x + 15 = 20 to solve. She says the solution is 1 2 . Which reason justifies her solution?

OpenStudy (anonymous):

does it say 12 or 1/2

OpenStudy (anonymous):

it should b 1/2

OpenStudy (anonymous):

yeah, if it is 1/2 than 10x +15 = 20 -15 -15 10x = 5 ÷10 ÷10 x = 1/2